OpenStudy (anonymous):

Help. I have a test on this today. Write an equation in slope intersect form from given info: (-1,2) // to y=1/2x-3. Then answer what is the m= and the //m=

OpenStudy (anonymous):

@rosimary Do you know slope intersect form?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

I think I figured out. 1 point is already given in slope int form. So the additional point I have to put in same way. So I use point slope form of y-y1=m(x-x1). So I y-2=1/2(x--1) then change to 2(y-2)2*1/2(x+1) then work out equation and leaves me with y=1/2x+5. Which means the //m=1/2 ... YES?

OpenStudy (anonymous):

m = 1/2 -> 2 = 1/2 (-1 ) + b => b = 2 + 1/2 = 5/2 Thus y = x/2 + 5/2

OpenStudy (anonymous):

That's right .... Thanks for catching that. Did I get the slope right?

OpenStudy (anonymous):

and the perpendicular one I just reverse slope and make into opposite whether (-) or (+) right?

OpenStudy (anonymous):

Great. Thank you. I think it was the way the question looked that it confused me.

OpenStudy (anonymous):

Observe the way I solving, there's no confusing with just 3 short lines !

OpenStudy (anonymous):

ok